OpenStudy (anonymous):

4Na+ZrCl4-> Zr +4NaCl Calculate the amount of zirconium produced (in kg) if a reactor were charged with 264.0 kg and 25.0 kg?

OpenStudy (anonymous):

I have worked it out but it seems the answer is wrong? 4Na = 91.96 g/mol 264000/91.96=2870.8134 moles of Na in 25Kg. Which is the limiting reagent? so therefore 2870.8134*91.22 = 261875.598/1000 = 261.875Kg of Zr produced no?